OpenStudy (please.help.me):

Use the sum-to-product formulas to find the exact value of the expression. 5 sin 75° + 5 sin 15°

OpenStudy (3mar):

May I help?

OpenStudy (please.help.me):

Yes, please!

OpenStudy (please.help.me):

Yes, please! @3mar

OpenStudy (please.help.me):

I plugged it into the formula but got the wrong answer

OpenStudy (3mar):

Of course. With my pleasure! So why you don't use angles (45) and (30) with their addition and subtraction to form the the angles (75) and (15)?

OpenStudy (please.help.me):

2*5 Sin ( (75°+15°)/2) cos( (75°-15°)/2)

OpenStudy (please.help.me):

10Sin (45°) Cos(30°)

OpenStudy (please.help.me):

That was my final answer but it was marked wrong

OpenStudy (3mar):

Why /2? \(2*5 Sin ( (75°+15°)\color{red}{/2}) cos( (75°-15°)\color{red}{/2})\)

OpenStudy (please.help.me):

Yes! Two is part of the formula so I divided it into the final answer

OpenStudy (3mar):

No sister! How did you get that\(10Sin (45°) Cos(30°)\) form this \(2*5 Sin ( (75°+15°)/2) cos( (75°-15°)/2)\)?

OpenStudy (3mar):

I think this would help you.

OpenStudy (please.help.me):

I did (75°+15°)/2= 45

jhonyy9 (jhonyy9):

@3mar nice job - hope will can understanding the asker what is the key how is possible solving this easy exercise

OpenStudy (please.help.me):

I am using the sum-to-product formula

jhonyy9 (jhonyy9):

@TheSmartOne please sorry i not can reply to you just in this form again too so i m here everyday to help and so in this way i ve arrived this 83"s level till today

jhonyy9 (jhonyy9):

but thank you for your question

OpenStudy (irishboy123):

\(5 \sin 75° + 5 \sin 15°\) \(= 5 ( \sin 75° + \sin 15°)\) for starters then from: \(\large \sin \theta \pm \sin \varphi =2\sin \left({\frac {\theta \pm \varphi }{2}}\right)\cos \left({\frac {\theta \mp \varphi }{2}}\right)\) \(5 \left( \sin 75^o + \sin 15^o =2\sin \left({\frac {75^o + 15^o }{2}}\right)\cos \left({\frac {75^o - 15^o }{2}}\right) \right)\)

OpenStudy (3mar):

So I mean: \[\large{ ~~~~5\sin(75)+5\sin(15)\\=5\sin(45+30)+5\sin(45-30)\\=5[\sin45\cos30+\cos30\sin45]+5[\sin45\cos30-\cos45\sin30]}\] Can you continue from here?

jhonyy9 (jhonyy9):

yes @3mar tis was my idea too - the best easy way to solve it

OpenStudy (please.help.me):

5[sin45cos30+cos30sin45]+5[sin45cos30−cos45sin30] 5[2sin45 2cos30]+5[sin45cos30−cos45sin30] ? Is this right?

OpenStudy (please.help.me):

But I was told to solve it with Product to Sum formulas http://www.sosmath.com/trig/prodform/prodform.html

jhonyy9 (jhonyy9):

thanky ou @3mar - sorry but i not can reply to you just in this way bc. my submitt buton is blocked again too

OpenStudy (please.help.me):

@IrishBoy123 That isn't the final answer is it?

OpenStudy (3mar):

So do you mean this?

OpenStudy (please.help.me):

I understand the last step from @IrishBoy123 And that is what I had in the beginning, but how to I simplify it for my final answer?

jhonyy9 (jhonyy9):

@3mar sorry but there where you wrote firstly the sin75 = sin(45+30) so please check it bc. there is one mistake sin45cos30 +cos30sin45 so in this way is wrong

jhonyy9 (jhonyy9):

So I mean: 5sin(75)+5sin(15)=5sin(45+30)+5sin(45−30)=5[sin45cos30+cos30sin45]+5[sin45cos30−cos45sin30] Can you continue from here?

jhonyy9 (jhonyy9):

do you see it now ?

OpenStudy (please.help.me):

@3mar No, I am using the Sum to Product Formulas

OpenStudy (3mar):

@jhonyy9 Thanks a lot. I apology, my mistake! \[\large{ ~~~~5\sin(75)+5\sin(15)\\=5\sin(45+30)+5\sin(45-30)\\=5[\sin45\cos30+\cos45\sin30]+5[\sin45\cos30-\cos45\sin30]}\] But you are right, @Please.help.me

OpenStudy (irishboy123):

\(5 \left( 2\sin \left({\frac {75^o + 15^o }{2}}\right)\cos \left({\frac {75^o - 15^o }{2}}\right) \right)\) \(= 10 \left( \sin {\frac {90^o }{2}} . \cos {\frac {60^o }{2}} \right)\) \(= 10 \left( \frac {1 }{\sqrt 2} \frac{\sqrt{3}}{2} \right) = 5 \sqrt {\frac { 3 }{2}} \)

OpenStudy (3mar):

So you need to do like this:  You simplify!